Table 1: Local Group Galaxies

Galaxy  Other Name  α_{2000}  δ_{2000}  l  b  Type  Subgroup Images Notes
(1)  (2)  (3)  (4)  (5)  (6)  (7)  (8)  (9)  (10)
WLM  DDO 221  00 01 58 −15 27.8 75.9 −73.6 IrrIV-V  LGC  1
NGC 55  00 15 08 −39 13.2 332.7 −75.7 IrrIV  LGC  3,4,51
IC 10  UGC 192  00 20 25 +59 17.5 119.0 −3.3 dIrr  M31  5,6
NGC 147  DDO 3  00 33 12 +48 30.5 119.8 −14.3 dSph/dE5  M31  7,51
And III  00 35 17 +36 30.5 119.3 −26.2 dSph  M31  8
NGC 185  UGC 396  00 38 58 +48 20.2 120.8 −14.5 dSph/dE3p  M31  7,51
NGC 205  M110  00 40 22 +41 41.4 120.7 −21.1 E5p/dSph-N  M31  7,51  E
M32  NGC 221  00 42 42 +40 51.9 121.2 −22.0 E2  M31  9,10,51
M31  NGC 224  00 42 44 +41 16.1 121.2 −21.6 SbI-II  M31  10,50,51
And I  00 45 43 +38 00.4 121.7 −24.9 dSph  M31  8
SMC  NGC 292  00 52 44 −72 49.7 302.8 −44.3 IrrIV-V  MW  11,12,51
Sculptor  01 00 09 −33 42.5 287.5 −83.2 dSph  MW  13,49
LGS 3  Pisces  01 03 53 +21 53.1 126.8 −40.9 dIrr/dSph  M31  14
IC 1613  DDO 8  01 04 54 +02 08.0 129.8 −60.6 IrrV  M31/LGC 15-17
And II  01 16 27 +33 25.7 128.9 −29.2 dSph  M31  8  A
M33  NGC 598  01 33 51 +30 39.6 133.6 −31.3 ScII-III  M31  17,18,51
Phoenix  01 51 06 −44 26.7 272.2 −68.9 dIrr/dSph  MW/LGC 19,20
Fornax  02 39 59 −34 27.0 237.1 −65.7 dSph  MW  21
EGB 0427+63 UGCA 92  04 32 01 +63 36.4 144.7 +10.5 dIrr  M31  23
LMC  05 23 34 −69 45.4 280.5 −32.9 IrrIII-IV  MW  11,12,51
Carina  06 41 37 −50 58.0 260.1 −22.2 dSph  MW  24
Leo A  DDO 69  09 59 24 +30 44.7 196.9 +52.4 dIrr  MW/N3109 29,51
Sextans B  DDO 70  10 00 00 +05 19.7 233.2 +43.8 dIrr  N3109  27,28,51
NGC 3109  DDO 236  10 03 07 −26 09.5 262.1 +23.1 IrrIV-V  N3109  29,30,51
Antlia  10 04 04 −27 19.8 263.1 +22.3 dIrr/dSph  N3109  31  B
Leo I  DDO 74  10 08 27 +12 18.5 226.0 +49.1 dSph  MW  32
Sextans A  DDO 75  10 11 06 −04 42.5 246.2 +39.9 dIrr  N3109  33,51
Sextans  10 13 03 −01 36.9 243.5 +42.3 dSph  MW  34
Leo II  DDO 93  11 13 29 +22 09.2 220.2 +67.2 dSph  MW  52
GR 8  DDO 155  12 58 40 +14 13.0 310.7 +77.0 dIrr  GR8  36,37,54
Ursa Minor  DDO 199  15 09 11 +67 12.9 105.0 +44.8 dSph  MW  38
Draco  DDO 208  17 20 19 +57 54.8 86.4 +34.7 dSph  MW  39
Milky Way  17 45 40 −29 00.5  0.0  0.0 Sbc  MW
Sagittarius  18 55 03 −30 28.7  5.6 −14.1 dSph-N  MW  40  E
SagDIG  UKS1927-177 19 29 59 −17 40.7 21.1 −16.3 dIrr  LGC  41,53
NGC 6822  DDO 209  19 44 56 −14 48.1 25.3 −18.4 IrrIV-V  LGC  42,43,51
DDO 210  Aquarius  20 46 46 −12 51.0 34.0 −31.3 dIrr/dSph  LGC  36,44,54  C
IC 5152  22 02 42 −51 17.7 343.9 −50.2 dIrr  LGC  51
Tucana  22 41 50 −64 25.2 322.9 −47.4 dSph  LGC  47,48  D
UKS2323-326 UGCA 438  23 26 27 −32 23.3 11.9 −70.9 dIrr  LGC  41
Pegasus  DDO 216  23 28 34 +14 44.8 94.8 −43.5 dIrr/dSph  LGC  25,44,51

EXPLANATION OF COLUMNS OF TABLE 1: Column 1: Galaxy name. Entries denoted in italics refer to the five giant Local Group galaxies that are not discussed in this review in any detail; Columns 2: A common alternative name; Columns 3 and 4: Right ascension and declination for epoch J2000.0, respectively; Columns 5 and 6: Galactic longitude and latitude, respectively; Column 7: Galaxy type following van den Bergh (1994a); Column 8: Subgroup membership within the Local Group; Column 9: References to optical and 21-cm images (if available); Column 10: Special notes.

NOTES FOR TABLE 1:

A. The position listed here is based on an independent measurement by Paul Hodge (private communication). B. Also known as PGC 29194 (Fouqué et al 1990) before it was re-discovered as a probable Local Group

member by Whiting et al (1997). A thorough discussion of ‘pre-discovery’ observations of Antlia is provided by Aparicio et al (1997a).

C. Marconi et al (1990) claimed that the original position of DDO 210 (Fisher and Tully 1975) was significantly in error. However, Lo et al (1993) noted that the original position seemed to be correct, and Marconi et al (1995) subsequently agreed. The position listed here corresponds to the original one from Fisher and Tully (1975).

D. Tucana was known prior to its re-discovery by Lavery and Mighell (1992) who thoroughly document all listings of the galaxy in earlier catalogs. Lavery and Mighell (1992) did first claim Tucana to be a possible Local Group member.

E. The ‘N’ suffix has been added to indicate that these systems may be nucleated dwarfs.

REFERENCES FOR TABLE 1: [1] Sandage & Carlson (1985b); [2] Lautsen et al (1977); [3] Hummel et al (1986); [4] Puche et al (1991); [5] de Vaucouleurs & Ables (1965); [6] Shostak & Skillman (1989); [7] Young & Lo (1997a); [8] Caldwell et al (1992); [9] Kent (1987); [10] Hodge (1981); [11] de Vaucouleurs & Freeman (1972); [12] Mathewson & Ford (1984); [13] Carignan et al (1998); [14] Young & Lo (1997b); [15] Ables (1971); [16] Lake & Skillman (1989); [17] Sandage (1961); [18] Corbelli et al (1989); [19] van de Rydt et al (1991); [20] Carignan et al (1991); [21] Hodge (1971); [22] Saha & Hoessel (1991); [23] Hoessel et al (1988); [24] Smecker-Hane et al (1994); [25] Sandage (1986b); [26] Young & Lo (1996a); [27] Sandage & Carlson (1985a); [28] Skillman et al (1988); [29] Sandage & Carlson (1988); [30] Jobin & Carignan (1990);

[31] Whiting et al (1997); [32] Hodge (1963a); [33] Sandage & Carlson (1982); [34] Mateo et al (1995a); [35] Deleted; [36] Fisher & Tully (1979); [37] Carignan et al (1990); [38] van Agt (1967); [39] Baade & Swope (1961); [40] Ibata et al (1997); [41] Longmore et al (1978); [42] Hodge (1978); [43] Gottesman & Weliachew (1977); [44] Lo et al (1993); [45] Deleted; [46] Deleted; [47] Lavery & Mighell (1992); [48] Oosterloo et al (1996); [49] van Agt (1978); [50] Hodge (1992a); [51] Sandage & Bedke (1994); [52] Vogt et al (1995); [53] Cesarsky et al (1977); [54] Hopp & Schulte-Ladbeck (1995).